1.     A truck moving at 35 m/s collide with a car moving at 50 m/s. After collision, both truck and car move together at the same speed. Mass of the truck is 1,500 kg and the mass of the car is 700 kg. What is the velocity of both truck and car after collision?



1
Expert's answer
2022-01-31T07:33:47-0500

Answer

Take truck direction as positive

Mass of truck

"m_t=1500kg \\
Speed \space v_t=35m/s"

Mass of car

"m_c=700kg \\
Speed \space v_c=-50m/s"


Using momemtum conservation

"Before" collision =after collision

"m_tv_t+m_cv_c=v(m_t+m_c)"

So after collision speed of car and truck

"\frac{m_tv_t+m_cv_c}{(m_t+m_c) }=v"

Putting all values

"v=\frac{1500\times 35-700\times50}{(1500+700) }\\=7.96m/s"